# Tru-POMDP: Task Planning Under Uncertainty via Tree of Hypotheses and Open-Ended POMDPs

### v1

@@ 1 Introduction
These tasks pose fundamental challenges in open household environments. First, human instructions are inherently ambiguous and often omit critical details (e.g., what specific items are needed for a party). Second, many objects are hidden from the robot’s initial view, and the set of relevant objects is unbounded and diverse (e.g., cups, snacks, decorations). As a result, the robot faces open-ended uncertainty over both world states and goals, leading to a boundlessly large belief space. Effective task planning in this setting requires reasoning over long horizons under severe partial observability, while tackling open-ended uncertainty.
@@
Recent work has started to leverage large language models (LLMs) to bring commonsense reasoning into robot planning. One line of work treats LLMs as planners, leveraging their implicit knowledge and ability to reflect over past failures yao2022-react ; shinn2023-reflexion ; huang2022inner_LLM-Monologue-Planners ; huang2022language_LLM-ZeroShot-Planners ; ahn2022can ; liang2023code . However, such approaches often rely on implicit, unverified hypotheses about the world, and can fail catastrophically when the underlying hypotheses are incorrect. Another line combines LLMs with symbolic or probabilistic planners, for example, by prompting LLMs to generate POMDP models for informing LLM planning or suggesting default actions for search  zhao2023large-LLM_MCTS ; Lee-PrimeTheSearch ; Sun-LLM_POP ; xiong-etal-2024-watchEveryStep ; curtis2025llmguidedprobabilisticprograminduction ; chen2025robohorizonllmassistedmultiviewworld . While more structured, these methods still commit to a single high-probability hypothesis and fail to model open-ended uncertainty, which limits robustness to ambiguity and occlusion.
@@
To overcome these limitations, we propose Tru-POMDP, a new algorithm for task planning under uncertainty that tightly integrates commonsense reasoning by LLMs with explicit belief tracking and principled POMDP planning. Tru-POMDP operates in three stages: (1) it constructs a tree of hypotheses over possible world states and task goals using hierarchical LLM queries, to form a commonsense belief; (2) it fuses the LLM-generated belief with Bayesian filtering to construct a hybrid belief that is both diverse and consistent during updates; and (3) it solves for an optimal policy using online belief tree search with a dynamically constructed action space and guided by an LLM-generated rollout policy. This design leverages the generalization power of LLMs while maintaining rigorous, verifiable reasoning over open-ended uncertainty.
@@
We evaluate Tru-POMDP on object rearrangement tasks across five diverse kitchen environments from RoboCasa nasiriany2024robocasa , involving open-ended instructions, ambiguous user intentions, and a large number of hidden spaces. Results show that Tru-POMDP effectively tackles challenges brought by open-ended uncertainties. Through the integration of LLM-based belief generation with POMDP planning, Tru-POMDP significantly outperforms both LLM agents and LLM-augmented tree search planners, producing higher success rates and plan qualities with lower LLM token usage. Tru-POMDP also benefits from structured belief modeling with LLMs and hybrid belief tracking that combines LLM belief with Bayesian updates.
@@
  • The first framework to integrate LLM-based reasoning with principled POMDP planning for household tasks.
@@
  • A POMDP model for open-ended object rearrangement tasks and a practical belief-tree search planner for solving such tasks efficiently under large-scale uncertainty.
@@ 2.1 Online POMDP Planning
Partially Observable Markov Decision Processes (POMDPs) provide a principled framework for decision-making under uncertainty, where the true state of the world is only partially and noisily observed. A POMDP is formally defined as a 7-tuple $(S,A,Z,T,O,R,b_{0})$, where $S$ is the set of states, $A$ is the set of actions, $Z$ is the set of observations, $T(s^{\prime}|s,a)$ is the transition model, $O(z|s^{\prime},a)$ is the observation model, $R(s,a)$ is the reward function, and $b_{0}$ is the initial belief over states. Online POMDP planning targets for single-query planning or even real-time planning, by computing policies via belief tree search. The agent maintains a belief—a probability distribution over possible world states—using Bayesian filtering. At each decision step, it constructs a belief tree by simulating future actions and observations NanYe2017DESPOT , and applies Bellman’s principle kaelbling1998planning in the tree to compute the best policy given the current belief. Only the immediate action is executed, and the process is repeated at each time step, allowing the planner to adapt to new observations in an online fashion (e.g., replan at 1 Hz).
@@ 2.2 Robot Planning with Large Language Models
Recent advances in large language models (LLMs) have enabled robots to handle tasks expressed in open-vocabulary language and to incorporate commonsense knowledge into the planning process. Broadly, two directions have emerged. The first uses LLMs as planners, either generating open-loop action sequences via prompt engineering or fine-tuning huang2022language_LLM-ZeroShot-Planners ; huang2022inner_LLM-Monologue-Planners ; singh2023progprompt ; chen2024agent ; ahn2022can ; liang2023code ; zhou2024isr , or producing closed-loop actions through iterative feedback and reflection shinn2023-reflexion ; yao2022-react ; Sun-AdaPlanner ; wang2024learning-NAT ; song2024-ETO ; yang2025selfgoal ; raman2022planning ; bhat2024grounding . While flexible, these approaches rely on implicit reasoning and often lack robustness when their internal hypotheses diverge from reality. The second direction combines LLMs with explicit planning algorithms. For example, some work uses LLMs to generate PDDL domains and problems, which are then solved by symbolic planners liu2023-LLMP ; han2024-interpret ; Birr2024-AUTOGPT_P ; dagan2023dynamic ; guan2023leveraging . Others combine LLMs with Monte Carlo Tree Search (MCTS), using the LLM to hypothesize initial states or initial solutions zhao2023large-LLM_MCTS ; Lee-PrimeTheSearch ; qiao2024-agent ; shi2025MC_DML . However, these methods generally assume full observability and deterministic goals, ignoring the inherent uncertainty in both human intention and world states. As a result, their solutions are often suboptimal in real-world deployments.
@@ 2.3 Planning under Uncertainty with Large Language Models
A few recent efforts have explored combining LLM reasoning with the structure of POMDPs. Some approaches use LLMs to generate elements of a POMDP model—such as beliefs or observation likelihoods—which are then fed back into the LLM to improve its planning ability under uncertainty Sun-LLM_POP ; xiong-etal-2024-watchEveryStep ; curtis2025llmguidedprobabilisticprograminduction ; chen2025robohorizonllmassistedmultiviewworld . However, these methods still rely heavily on the LLM’s implicit reasoning and lack the formal guarantees provided by explicit belief tracking and tree-search planning. Other work uses the LLM to generate a single most likely hypothesis about goals and world states, and then applies MCTS for planning zhao2023large-LLM_MCTS ; hazra2024saycanpay ; ren2023robotsaskhelpuncertainty . This strategy fails when the true goal diverges from the most likely one, and it incurs significant computational cost due to repeated LLM queries within the tree search he2024wordsactionsunveilingtheoretical . In addition, MCTS often requires a known set of objects, restricting it to a closed domain. In contrast, our approach uses the LLM to construct a tree of plausible hypotheses representing a diverse and open-ended belief over hidden states and goals. We combine this with Bayesian filtering to maintain a consistently updatable belief. Finally, we apply belief tree search over this hybrid belief with a dynamically constructed action space to compute near-optimal policies under open-ended uncertainty. This integration allows us to leverage the generalization and commonsense reasoning of LLMs while retaining the robustness and rigor of POMDP planning.
@@ 4.2 Hybrid Belief Update
The Hybrid Belief Update module maintains a consistent and efficient particle belief over partially observed world states and placement goals. It fulfills two objectives: (1) ensuring smooth and reliable belief updates through Bayesian filtering, and (2) reducing the frequency of expensive LLM queries by only invoking them when necessary. The algorithm monitors the total weight of the particle belief. If the belief retains sufficient weight—indicating strong agreement with the current observation—it is updated using a particle filter. If the weight falls below a threshold (e.g., $0.31-\epsilon=$ 0.3), new particles are generated by invoking the Tree of Hypotheses module (Section 4.1) to replenish the belief.
@@ 4.3 Online POMDP Planning
This module plans the optimal policy given the current belief over uncertain world states and goals. Our planner is built upon DESPOT NanYe2017DESPOT , a sampling-based online POMDP algorithm that provides asymptotic optimality guarantees. We extend DESPOT with two key innovations: (1) a dynamic action space constructed from belief particles, and (2) an LLM-generated rollout policy that injects commonsense domain knowledge while maintaining computational efficiency. Tru-POMDP operates in an online planning setting: at each time step, it executes the optimal action for the current belief and replans after receiving new observations.
@@
Belief Tree Search. We adapt the DESPOT algorithm to recursively explore the space of future action-observation sequences from the current belief. At each time step, a sparse belief tree is constructed: nodes represent future belief states (particle beliefs), and branches represent possible actions and the resulting observations, simulated using the POMDP transition and observation models. Each belief node also maintains a value estimate, representing the expected cumulative reward that can be achieved under optimal behavior conditioned on that belief. These values are updated iteratively using Bellman’s principle NanYe2017DESPOT . The planner leverages value estimates as tree search heuristics: it selects actions that maximize the value estimate, and expands observation branches based on excess uncertainty. This ensures targeted exploration of promising belief trajectories, ensuring efficient convergence of values and allowing planning to terminate at any time and output high-quality decisions.
LLM-Generated Rollout Policy. Following the DESPOT framework, leaf nodes in our belief tree require rolling-out a default policy to provide a heuristic lower bound on their value. However, unlike prior work that simply uses a random rollout policy zhao2023large-LLM_MCTS , we query the LLM to synthesize a rollout policy in code. The LLM is provided with a domain description that includes the task objective, action definitions, state abstraction, and constraints—but omits any environment-specific information. The resulting policy generalizes across tasks in the same domain and is reused throughout the search. During rollouts, this policy is simulated for each rolled-out belief to estimate cumulative rewards from the leaves.
@@ 5.1 Experimental Setup
We evaluate Tru-POMDP in five kitchen environments from RoboCasa nasiriany2024robocasa , which highlights a large number of hidden areas: One Wall, One Wall with Island, L-Shaped, L-Shaped with Island, and Galley. Each scene contains up to 40 AREA nodes with semantically-rich names, of which up to 29 are initially closed. Tasks are procedurally generated using an LLM-assisted pipeline, emphasizing vague task instructions with substantial implicit information and ambiguous object references. For each task, we populate up to 20 objects (including distractors) in the scene using commonsense priors, and define a goal set $\mathcal{P}$ containing various plausible target object combinations and goal placements grounded in the scene. This setup ensures each task is both solvable and exhibits ambiguity in human intention. Example scene with instruction and goal sets are listed in Appendix A.3 and A.4.
@@
Experiments are run on a local machine equipped with a 12th Gen Intel® Core™ i7-12700KF CPU (20 threads), without GPU acceleration. All methods consistently use GPT-4.1 as the LLM.
@@ 5.2 Comparison Results
  • ReAct yao2022-react : A closed-loop LLM-based planner that selects actions based on current observations and immediate feedback from the environment.
  • Reflexion shinn2023-reflexion : An extension of ReAct that introduces a reflection module. When repeated failures are encountered, it analyzes the execution history and generates a revised plan to guide subsequent decisions. For fair comparison in an online planning setup, we disable environment resets and trigger reflection after three consecutive failed actions.
@@
  • LLM-MCTS zhao2023large-LLM_MCTS : A tree search method that uses an LLM to generate a single maximum-likelihood hypothesis of the task goal and hidden object placements. It then performs Monte Carlo Tree Search (MCTS), calling the LLM repeatedly to guide the action selection in simulation procedure of MCTS.
@@
Tru-POMDP vs. Tree Search planners. LLM-MCTS struggles with planning efficiency due to frequent LLM calls during simulation, exceeding the 10-minute time limit in all tasks. Tru-POMDP mitigates this using offline LLM-generated policy, expressed as C++ code. Tru-POMDP also integrates LLM reasoning with rigorous belief-space planning and performs tree search over a structured particle belief, enabling it to solve ambiguous tasks more effectively than all baselines. Note that Tru-POMDP achieved this using minimum consumption of LLM tokens (Figure 3). See Appendix A.7 for visualization of the planned results.
@@ 6 Conclusion and Limitations
Nevertheless, our approach currently has several limitations. The Tree of Hypotheses incurs computational overhead due to multiple LLM calls, which could be mitigated by fine-tuning an LLM for faster belief generation. Our experiments also assumed noise-free perception and deterministic actions. However, the general POMDP formulation naturally accommodates perception noise and stochastic outcomes by adding appropriate probabilistic models. Finally, our evaluation focused on object rearrangement tasks. Extending the approach to larger action spaces and reasoning over additional unknown object attributes remains an important direction for future work.

### v2

# arXiv:2506.02860v2[cs.RO] 01 Mar 2026

@@ 1 Introduction
While these challenges can be positioned within the framework of planning under uncertainty—where a robot must hedge against imperfect perception, unobservable environment states, and stochastic action outcomes—the difficulties in open household environments extend far beyond standard formulations. First, human instructions are inherently ambiguous and often omit crucial details (e.g., what specific items are needed for a party). Second, many objects are hidden from the robot’s initial view, and the set of relevant objects is unbounded and diverse (e.g., cups, snacks, decorations). As a result, the robot must reason over open-ended uncertainty, where fundamental components of the planning problem—such as the state space (e.g., what objects exist), the action space (e.g., what objects to operate on) and task goals (e.g., what the human want)—cannot be predefined. This setting leads to an effectively boundless belief space, making long-horizon planning extremely challenging.
@@
To address open-world problems, recent work has started to leverage large language models (LLMs) to bring commonsense reasoning into robot planning. One line of work treats LLMs as planners, leveraging their implicit knowledge and ability to reflect over past failures Yao et al. (2022); Shinn et al. (2023); Huang et al. (2022b, a); Ahn et al. (2022); Liang et al. (2023). However, such approaches often rely on implicit, unverified hypotheses about the world, and can fail catastrophically when the underlying hypotheses are incorrect. Another line combines LLMs with symbolic or probabilistic planners, for example, by prompting LLMs to generate POMDP models for informing LLM planning or suggesting default actions for search  Zhao et al. (2023); Lee et al. (2024); Sun et al. (2024); Xiong et al. (2024); Curtis et al. (2025); Chen et al. (2025). While more structured, these methods still commit to a single high-probability hypothesis or operate in a closed domain, which limits robustness in the open-world.
@@
To overcome these limitations, we propose Tru-POMDP, a task planner built upon an open-ended POMDP formulation, that tightly integrates commonsense reasoning by LLMs with explicit belief tracking and principled POMDP planning. Tru-POMDP operates in three stages: (1) it constructs a tree of hypotheses over possible world states and task goals using hierarchical LLM queries, to form a commonsense belief; (2) it fuses the LLM-generated belief with Bayesian filtering to construct a hybrid belief that is both diverse and consistent during updates; and (3) it solves for an optimal policy using online belief tree search with a dynamically constructed action space and guided by an LLM-generated rollout policy. This design leverages the generalization power of LLMs while maintaining rigorous, verifiable reasoning over open-ended uncertainty.
@@
We evaluate Tru-POMDP on object rearrangement tasks across five diverse kitchen environments from RoboCasa Nasiriany et al. (2024), involving open-ended instructions, ambiguous user intentions, and a large number of hidden spaces. Results show that Tru-POMDP effectively tackles challenges brought by open-ended uncertainties. Through the integration of LLM-based belief generation with POMDP planning, Tru-POMDP significantly outperforms both LLM agents and LLM-augmented tree search planners, producing higher success rates and plan qualities with lower LLM token usage. Tru-POMDP also benefits from structured belief modeling with LLMs and hybrid belief tracking that combines LLM belief with Bayesian updates.
@@
  • The first framework to address “open-ended POMDPs”, integrating LLM-based reasoning with principled POMDP planning for household tasks.
@@
  • A POMDP model for open-world object rearrangement tasks and a practical belief-tree search planner for solving such tasks efficiently under large-scale uncertainty.
@@ 2.1 Online POMDP Planning
Partially Observable Markov Decision Processes (POMDPs) provide a principled framework for decision-making under uncertainty, where the true state of the world is only partially and noisily observed. A POMDP is formally defined as a 7-tuple $(S,A,Z,T,O,R,b_{0})$, where $S$ is the set of states, $A$ is the set of actions, $Z$ is the set of observations, $T(s^{\prime}|s,a)$ is the transition model, $O(z|s^{\prime},a)$ is the observation model, $R(s,a)$ is the reward function, and $b_{0}$ is the initial belief over states. Online POMDP planning targets for single-query planning or even real-time planning, by computing policies via belief tree search. The agent maintains a belief—a probability distribution over possible world states—using Bayesian filtering. At each decision step, it constructs a belief tree by simulating future actions and observations Ye et al. (2017), and applies Bellman’s principle Kaelbling et al. (1998) in the tree to compute the best policy given the current belief. Only the immediate action is executed, and the process is repeated at each time step, allowing the planner to adapt to new observations in an online fashion (e.g., replan at 1 Hz). A POMDP model requires a full specification of all possible states, actions, and observations, which limits its use to "closed-domain" problems.
@@ 2.2 Robot Planning with Large Language Models
Recent advances in large language models (LLMs) have enabled robots to handle tasks expressed in open-vocabulary language and to incorporate commonsense knowledge into the planning process. Broadly, two directions have emerged. The first uses LLMs as planners, either generating open-loop action sequences via prompt engineering or fine-tuning Huang et al. (2022a, b); Singh et al. (2023); Chen et al. (2024); Ahn et al. (2022); Liang et al. (2023); Zhou et al. (2024), or producing closed-loop actions through iterative feedback and reflection Shinn et al. (2023); Yao et al. (2022); Sun et al. (2023); Wang et al. (2024); Song et al. (2024); Yang et al. (2025); Raman et al. (2022); Bhat et al. (2024). While flexible, these approaches rely on implicit reasoning and often lack robustness when their internal hypotheses diverge from reality. The second direction combines LLMs with explicit planning algorithms. For example, some work uses LLMs to generate PDDL domains and problems, which are then solved by symbolic planners Liu et al. (2023); Han et al. (2024); Birr et al. (2024); Dagan et al. (2023); Guan et al. (2023). Others combine LLMs with Monte Carlo Tree Search (MCTS), using the LLM to hypothesize initial states or initial solutions Zhao et al. (2023); Lee et al. (2024); Qiao et al. (2024); Shi et al. . However, these methods generally assume full observability and deterministic goals, ignoring the inherent uncertainty in both human intention and world states. As a result, their solutions are often suboptimal in real-world deployments.
@@ 2.3 Planning under Uncertainty with Large Language Models
A few recent efforts have explored combining LLM reasoning with the structure of POMDPs. Some approaches use LLMs to generate elements of a POMDP model—such as beliefs or observation likelihoods—which are then fed back into the LLM to improve its planning ability under uncertainty Sun et al. (2024); Xiong et al. (2024); Curtis et al. (2025); Chen et al. (2025). However, these methods still rely heavily on the LLM’s implicit reasoning and lack the formal guarantees provided by explicit belief tracking and tree-search planning. Other work combines LLMs with tree search. One possibility is to generate a single most likely hypothesis about goals and world states, and then applies MCTS for planning Hazra et al. (2024); Ren et al. (2023). This strategy fails when the true goal diverges from the most likely one. LLM-MCTS Zhao et al. (2023) proposed to plan with LLM-generated state distributions. However, it uses a "flat" belief that only considers a single aspect of uncertainty on the initial placement of objects. It operates in a closed domain, assuming the object set is known beforehand, and incurs significant computational cost due to repeated LLM queries within the tree search He et al. (2024). In contrast, Tru-POMDP offers a scalable and flexible approach to address open-ended uncertainty. It uses the LLM to construct a tree of plausible hypotheses representing a complex, multi-aspect, and open-ended belief over entirely unknown objects and ambiguous human goals. We combine this with Bayesian filtering to maintain a consistently updatable belief. Finally, we apply belief tree search over this hybrid belief, using a dynamically constructed action space and a pre-compiled, LLM-generated rollout policy to efficiently guide optimal search.
@@ 4.2 Hybrid Belief Update
The Hybrid Belief Update module maintains a consistent and efficient particle belief over partially observed world states and placement goals. It fulfills two objectives: (1) ensuring smooth and reliable belief updates through Bayesian filtering, and (2) reducing the frequency of expensive LLM queries by only invoking them when necessary. The algorithm monitors the total weight of the particle belief. If the belief retains sufficient weight—indicating strong agreement with the current observation—it is updated using a particle filter. If the weight falls below a threshold (e.g., $1-\epsilon=0.3$), new particles are generated by invoking the Tree of Hypotheses module (Section 4.1) to replenish the belief.
@@ 4.3 Online POMDP Planning
This module plans the optimal policy given the current belief over uncertain world states and goals. Our planner is built upon DESPOT Ye et al. (2017), a sampling-based online POMDP algorithm that provides asymptotic optimality guarantees. We extend DESPOT with two key innovations: (1) a dynamic action space constructed from belief particles, and (2) an LLM-generated rollout policy that injects commonsense domain knowledge while maintaining computational efficiency. Tru-POMDP operates in an online planning setting: at each time step, it executes the optimal action for the current belief and replans after receiving new observations.
@@
Belief Tree Search. We adapt the DESPOT algorithm to recursively explore the space of future action-observation sequences from the current belief. At each time step, a sparse belief tree is constructed: nodes represent future belief states (particle beliefs), and branches represent possible actions and the resulting observations, simulated using the POMDP transition and observation models. Each belief node also maintains a value estimate, representing the expected cumulative reward that can be achieved under optimal behavior conditioned on that belief. These values are updated iteratively using Bellman’s principle Ye et al. (2017). The planner leverages value estimates as tree search heuristics: it selects actions that maximize the value estimate, and expands observation branches based on excess uncertainty. This ensures targeted exploration of promising belief trajectories, ensuring efficient convergence of values and allowing planning to terminate at any time and output high-quality decisions.
LLM-Generated Rollout Policy. Following the DESPOT framework, leaf nodes in our belief tree require rolling-out a default policy to provide a heuristic lower bound on their value. However, unlike prior work that simply uses a random rollout policy Zhao et al. (2023), we query the LLM to synthesize a rollout policy in code. The rollout policy is a mapping from the current state $s=(G_{s},p_{s})$ to an action $a$, where $G_{s}$ is the scene graph representing the hypothesized world state, and $p_{s}$ is the placement goal. This mapping is implemented as a C++ function generated by an LLM. The generator LLM receives a general language description of the object rearrangement domain but is given no information about any specific task instance. The resulting policy generalizes across tasks in the same domain and is reused throughout the search. During rollouts, this policy is simulated for each rolled-out node to estimate cumulative rewards from the leaves. See Appendix A.2 for the generator prompt and the code of the rollout policy.
@@ 5.1 Experimental Setup
We evaluate Tru-POMDP in five kitchen environments from RoboCasa Nasiriany et al. (2024), which highlights a large number of hidden areas: One Wall, One Wall with Island, L-Shaped, L-Shaped with Island, and Galley. Each scene contains up to 40 AREA nodes with semantically-rich names, of which up to 29 are initially closed. Tasks are procedurally generated using an LLM-assisted pipeline, emphasizing vague task instructions with substantial implicit information and ambiguous object references. For each task, we populate up to 20 objects (including distractors) in the scene using commonsense priors, and define a goal set $\mathcal{P}$ containing various plausible target object combinations and goal placements grounded in the scene. This setup ensures each task is both solvable and exhibits ambiguity in human intention. Example scene with instruction and goal sets are listed in Appendix A.3 and A.4.
@@
Experiments are run on a local machine equipped with a 12th Gen Intel® Core™ i7-12700KF CPU (20 threads), without GPU acceleration. All methods consistently use GPT‑4.1 as the LLM.
@@ 5.2 Comparison Results
  • ReAct Yao et al. (2022): A closed-loop LLM-based planner that selects actions based on current observations and immediate feedback from the environment.
  • Reflexion Shinn et al. (2023): An extension of ReAct that introduces a reflection module. When repeated failures are encountered, it analyzes the execution history and generates a revised plan to guide subsequent decisions. For fair comparison in an online planning setup, we disable environment resets and trigger reflection after three consecutive failed actions.
@@
  • LLM-MCTS Zhao et al. (2023): LLM augmented closed-domain POMDP planning. Given a known object set, LLM generates object-location probability vectors as belief. It then performs Monte Carlo Tree Search (MCTS), calling the LLM repeatedly to guide the action selection in simulation procedure of MCTS.
@@
Tru-POMDP vs. Tree Search planners. LLM-MCTS struggles with planning efficiency due to frequent LLM calls during simulation, exceeding the 10-minute time limit in all tasks. Tru-POMDP mitigates this using offline LLM-generated policy, expressed as C++ code. LLM-MCTS also operates on a closed object set, and can never accomplish a task when the true targets fall outside of the set. Tru-POMDP addresses this by constructing open-ended beliefs and dynamic action spaces. Tru-POMDP also integrates LLM reasoning with rigorous belief-space planning and performs tree search over a structured particle belief, enabling it to solve ambiguous tasks more effectively than all baselines. Note that Tru-POMDP achieved this using minimum consumption of LLM tokens (Figure 3). See Appendix A.7 for visualization of the planned results.
@@ 6 Conclusion and Limitations
Nevertheless, our approach currently has several limitations. The Tree of Hypotheses incurs computational overhead due to multiple LLM calls, which could be mitigated by fine-tuning an LLM for faster belief generation. Furthermore, Tru-POMDP leverages conditional independence between the placement of different objects to achieve scalable belief modeling, as this property naively supports factorization. When more complex dependency exists, systematic factorization is required to determine the belief structure. Next, our experiments also assumed noise-free perception and deterministic actions. However, the general POMDP formulation naturally accommodates perception noise and stochastic outcomes by adding appropriate probabilistic models. Finally, our evaluation focused on object rearrangement tasks. Extending the approach to larger action spaces and reasoning over additional unknown object attributes remains an important direction for future work.